# VP-MEL: Visual Prompts Guided Multimodal Entity Linking

### v3

@@ 7 Conclusion
In this paper, we introduce a new task, VP-MEL, aiming to link visual regions in image-text pairs to the corresponding entities in knowledge base, guided by visual prompts. To support this task, we construct a high-quality dataset, VPWiki, leveraging an automated annotation pipeline to enhance the efficiency of data annotation. To address the VP-MEL task, we propose FBMEL, a framework that can effectively utilize visual prompts to extract enhanced local visual features and generate supplementary textual information. FBMEL maintains a balance between visual and textual features, preventing excessive reliance on a single modality. Extensive experimental results demonstrate that FBMEL surpasses state-of-the-art methods. Moreover, VP-MEL significantly alleviates the constraints of mention words and expands the applicability of MEL to real-world scenarios.
@@ A.1 Experimental Settings
For our proposed model framework, we use pre-trained ViT-B/32 (Dosovitskiy et al., 2021) as the visual encoder, initialized with weights from CLIP-ViT-Base-Patch32111https://huggingface.co/openai/clip-vit-base-patch32, with $d_{v}$ and $d_{c}$ set to 96. The batch size is set to 128. For the text encoder, we select pre-trained BERT model (Devlin et al., 2019), setting the maximum input length for text to 40 and the output feature dimension $d_{t}$ to 512. We train and test on a device equipped with 4 Intel(R) Xeon(R) Platinum 8380 CPUs and 8 NVIDIA A800-SXM4-80GB GPUs.
@@ A.5 Feature Interaction Formula
A.6 Pipeline Module Settings
@@
WIKIDiverse WikiMEL
Sentences 7,405 22,070
M. in train 11,351 18,092
M. in valid 1,664 2,585
M. in test 2,078 5,169
Entities 132,460 109,976
Table 7: Statistics of WIKIDiverse and WikiMEL. M. denotes Mentions.
@@ A.7 WikiDiverse and WikiMEL
[…]
A.8 Additional Experiments and Experimental Descriptions
The knowledge base (KB) used in the WikiMEL and RichpediaMEL datasets differs from that of VPWiki and WIKIDiverse. To comprehensively assess the performance of the FBMEL framework in the MEL task, we retrain FBMEL on the WikiMEL and RichpediaMEL datasets. Importantly, neither visual prompts nor the Detective-VLM are employed during the retraining process. The experimental results are shown in Table 8. FBMEL is second only to MIMIC and MELOV, demonstrating its competitive performance as an MEL method.
@@ A.8 Additional Experiments and Experimental Descriptions
Figure 6: Case study for VP-MEL. Each row is a case, which contains Input, ground truth entity, and top three retrieved entities of three methods, i.e., FBMEL (ours), LLaVA-1.5 (Liu et al., 2024), MIMIC (Luo et al., 2023). Each retrieved entity is described by its Wikidata QID and entity name, with the entity marked with a checkmark indicating the correct one.
A.9 Case Study
To clearly demonstrate the proposed VP-MEL task and the FBMEL model, we conduct case studies and compare them against two strong competitors (i.e., LLaVA-1.5 and MIMIC), in Figure 6. As shown in Figure 6(a), in the first case, all three methods correctly predicted the entity. FBMEL makes full use of both image and text information, allowing it to more effectively distinguish between the different individuals in the image. LLaVA-1.5 may be overwhelmed by the textual information, while MIMIC struggles to identify the correct entity when the mention words are unavailable. In the second case, both LLaVA-1.5 and MIMIC retrieve Endeavour as the first choice. Only FBMEL, with the guidance of Visual Prompts and integration of textual information, correctly predicts the right entity. In Figure 6(b), we present the failed predictions. In the first case, when the content of the image interferes with the visual prompt, it impairs the reasoning process of FBMEL. The red box in the image bears a high similarity to the visual prompt. As a result, FBMEL incorrectly focuses on the wrong region of the image, ranking Donald Trump first. When FBMEL encounters difficulties in distinguishing the objects within the visual prompts, it leads to incorrect inferences. For example, in the second case, the distinguishing features of the two individuals in the image are obstructed, which causes FBMEL to struggle in differentiating between them. The image content in real-world data is often complex, which makes VP-MEL a challenging task. We hope that this task can be further refined and developed over time.

### v4

@@ 7 Conclusion
In this paper, we introduce a new task, VP-MEL, aiming to link visual regions in image-text pairs to the corresponding entities in knowledge base, guided by visual prompts. To support this task, we construct a high-quality dataset, VPWiki, leveraging an automated annotation pipeline to enhance the efficiency of data annotation. To address the VP-MEL task, we propose IIER, a framework that can effectively utilize visual prompts to extract enhanced local visual features and generate supplementary textual information. IIER maintains a balance between visual and textual features, preventing excessive reliance on a single modality. Extensive experimental results demonstrate that IIER surpasses state-of-the-art methods. Moreover, VP-MEL significantly alleviates the constraints of mention words and expands the applicability of MEL to real-world scenarios.
@@ A.1 Experimental Settings
For our proposed model framework, we use pre-trained ViT-B/32 (Dosovitskiy et al., 2021) as the visual encoder, initialized with weights from CLIP-ViT-Base-Patch32111https://huggingface.co/openai/clip-vit-base-patch32, with $d_{v}$ and $d_{c}$ set to 96. The number of epochs is set to 20, and the learning rate is tuned to $1\times 10^{-5}$. The batch size is set to 128. In the loss function, $\lambda$ is set to 1. For the text encoder, we select pre-trained BERT model (Devlin et al., 2019), setting the maximum input length for text to 40 and the output feature dimension $d_{t}$ to 512. Without including the VLM, the size of the trainable parameters is 153 M, and the total estimated model parameters size is 613 M. We train and test on a device equipped with 4 Intel(R) Xeon(R) Platinum 8380 CPUs and 8 NVIDIA A800-SXM4-80GB GPUs.
@@ A.5 Feature Interaction Formula
A.6 Data Annotation Pipeline
Please note that the pipeline serves as a preprocessing stage for data annotation. We use the Visual Entailment Module and the Visual Grounding Module to automatically annotate visual prompts in the images. While the accuracy of the pipeline is limited—such as its difficulty in distinguishing between specific individuals when multiple people are present—it still plays a crucial role in improving annotation efficiency. Due to these limitations, manual verification and re-annotation are necessary after pipeline processing. However, for annotators, making a simple "yes or no" judgment is much easier than selecting a specific individual. As a result, even with limited accuracy, the pipeline significantly boosts the overall efficiency of the annotation process.
@@ A.6 Data Annotation Pipeline
WIKIDiverse WikiMEL RichpediaMEL
Sentences 7,405 22,070 17,724
M. in train 11,351 18,092 12,463
M. in valid 1,664 2,585 1,780
M. in test 2,078 5,169 3,562
Entities 132,460 109,976 160,935
Table 7: Statistics of WIKIDiverse, WikiMEL, and RichpediaMEL. M. denotes Mentions.
@@ A.7 WikiDiverse and WikiMEL
IIER 88.93 95.69 96.73 84.63 93.27 95.30
Table 8: Baseline results marked with "*" according to Sui et al. (2024). We run each method three times with different random seeds and report the mean value of every metric. The best score is highlighted in bold.
A.8 Data Samples
We provide additional data samples categorized by entity type. The specific details can be found in Figure 7.
A.9 Additional Experiments
To comprehensively assess the performance of the IIER framework in the MEL task, we test IIER on the WikiMEL and RichpediaMEL datasets (Wang et al., 2022a). The statistics of WikiMEL and RichpediaMEL are shown in Table 7. Experimental results are shown in Table 8. The experimental results demonstrate that IIER remains highly competitive with state-of-the-art MEL method.
It is noted that within these two datasets, certain metrics of IIER exhibit values that are comparable to those of MELOV, such as H@1 and H@3 in the WikiMEL dataset. This may be attributed to the higher image quality and the homogeneous entity types (primarily $Person$) in WikiMEL and RichpediaMEL. When datasets contain fewer entity types and minimal image noise, the auxiliary information generated by IIER contributes less to performance improvement.
Nevertheless, IIER achieves the best performance on WikiDiverse, which includes a wider variety of entity types, and achieves a new SOTA for the VP-MEL task. As MEL increasingly addresses more complex scenarios, IIER shows significant potential for future advancements.
@@ A.9 Additional Experiments
Figure 6: Case study for VP-MEL. Each row is a case, which contains Input, ground truth entity, and top three retrieved entities of three methods, i.e., IIER (ours), LLaVA-1.5 (Liu et al., 2024), MIMIC (Luo et al., 2023). Each retrieved entity is described by its Wikidata QID and entity name, with the entity marked with a checkmark indicating the correct one.
Figure 7: Examples of the VPWiki dataset. Each row represents a sample corresponding to a specific entity type, which contains the entity type, image of mention, text of mention, image of entity, and entity.
A.10 Case Study
To clearly demonstrate the proposed VP-MEL task and the IIER model, we conduct case studies and compare them against two strong competitors (i.e., LLaVA-1.5 and MIMIC), in Figure 6. As shown in Figure 6(a), in the first case, all three methods correctly predicted the entity. IIER makes full use of both image and text information, allowing it to more effectively distinguish between the different individuals in the image. LLaVA-1.5 may be overwhelmed by the textual information, while MIMIC struggles to identify the correct entity when the mention words are unavailable. In the second case, both LLaVA-1.5 and MIMIC retrieve Endeavour as the first choice. Only IIER, with the guidance of Visual Prompts and integration of textual information, correctly predicts the right entity. In Figure 6(b), we present the failed predictions. In the first case, when the content of the image interferes with the visual prompt, it impairs the reasoning process of IIER. The red box in the image bears a high similarity to the visual prompt. As a result, IIER incorrectly focuses on the wrong region of the image, ranking Donald Trump first. When IIER encounters difficulties in distinguishing the objects within the visual prompts, it leads to incorrect inferences. For example, in the second case, the distinguishing features of the two individuals in the image are obstructed, which causes IIER to struggle in differentiating between them. The image content in real-world data is often complex, which makes VP-MEL a challenging task. We hope that this task can be further refined and developed over time.